A 1/A =B 1/B =C 1/C A BC +B AC +C AB =729 FIND A 1/A

A1/A=B1/B=C1/C
ABC+BAC+CAB=729
FIND A1/A

Grade:12th pass

2 Answers

Anoopam Mishra
8 years ago
A^{1/A} = B^{1/B} = C^{1/C}
=> A^{BC} = B^{CA} = C^{AB} = K (any\ variable)
=> 3K = 729
=> K = 243
=> => A^{1/A} = 243
 
Sparsh Chawla
5 years ago
A1/A = B1/B = C1/C = K 
A = KA; B = KB; C = KC 
ABC+BAC+CAB=729
KABC+KABC+KABC=729
3KABC=729
KABC=243
K= (243)1/ABC
A1/A = (243)1/ABC